(5,4) (0,2)
find the slope
state the x and y intercepts
put in to standard form

how do i find the x and y intercept

Answer by stanbon(48558)

The slope is (4-2)/(5-0)= 2/5
The form of the line is y=mx+b
You have m=2/5, x=0, y=2
Solve for "b" bu substituting those values:
2=(2/5)(0)+b
b = 2
The equation is y=(2/5)x+2
To get the x-intercept, let y=0
Then (2/5)x+2=0
2x+10=0
x=-5 (the x-intercept is (-5,0))
To get the y-intercept, let x=0
Then y=(2/5)(0)+2
y=2 (the y-intercept is (0,2))
Cheers,
Stan H.